# The foot, the fan, and the cuprate phase diagram: Fermi-volume-changing quantum phase transitions

### v6

@@ 1 Jan Zaanen and the cuprates
The other direction works directly with the SYK model, and modifies it to a more realistic representation of the underlying lattice scale physics, as reviewed in Ref. [20]. It is this direction which we will discuss in the rest of this article. A fruitful formulation, which has been the focus of recent work, has been the mapping of theories of quantum phase transitions in disordered metals to the two-dimensional complex Yukawa-Sachdev-Ye-Kitaev (2dYSYK) model [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 26, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], which we will discuss in Sections 3 and 4. With the benefit of hindsight, we can now see some of this progress could have been made after the early papers [15, 19, 16, 18] on what was then sometimes called the SY model. But the intervening developments on the connections between SYK, black holes, and holography were important in pointing to the correct direction, and Jan played a significant role in these. A possible holographic dual of the 2dYSYK model remains an interesting open question for future work, which we will not address here.
@@ 4 The ‘fan’: QPT without symmetry breaking
Instead the observations of Robarts et al. [109] are naturally interpreted as spinons. Note the evolution of their observations with increasing doping from a sharp spin-wave spectrum in Figs. 8a,b to a broad continuum in Figs. 8c-f. This matches Fig. 9, which shows the evolution of the dynamic spin structure factor of the insulating square lattice antiferromagnet with increasing second neighbor exchange $J_{2}$, as computed by Ferrari and Becca [113] using fermionic spinons (see also the connection of spinons in the unfrustrated square lattice to the quasi-one-dimensional limit [114, 115] . There is also a correspondence in the energy scale, measured in units of the first neighbor exchange $J_{1}\approx 130$ meV [116]. Indeed, the cuprate observations are the analog of those in the triangular lattice antiferromagnet KYbSe2 [61]. As discussed in Section 2 and Fig. 5, KYbSe2 displays a continuum at high energies which has been modeled in a theory of bosonic spinons [62, 61], along with weak antiferromagnetic order at low temperatures.
@@ 5 Confinement crossover
It is worth mentioning here an important low $T$ difference between FL* and the holon metal which could be significant for future transport experiments. As the holons are spinless, the area enclosed by a holon pocket is twice that of the hole pocket in FL* [144], and the holon pocket area is the same as that of the hole pocket in the ordered antiferromagnet.
@@
The confinement crossover from the pseudogap metal to the stripy SDW state at low temperatures remains to be understood [154]. In the fermionic spinon approach, it involves a delicate interplay between the charge order associated with the condensation of $B$, and the confinement of the fermionic spinons via the SU(2) gauge field [139]. The bosonic spinon approach [73, 74, 75, 142] may be better suited to understand this crossover, and insights from numerical studies [155] should also be helpful.

### v7

@@ 1 Jan Zaanen and the cuprates
The other direction works directly with the SYK model, and modifies it to a more realistic representation of the underlying lattice scale physics, as reviewed in Ref. [20]. It is this direction which we will discuss in the rest of this article. A fruitful formulation, which has been the focus of recent work, has been the mapping of theories of quantum phase transitions in disordered metals to the two-dimensional complex Yukawa-Sachdev-Ye-Kitaev (2dYSYK) model [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 26, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59], which we will discuss in Sections 3 and 4. With the benefit of hindsight, we can now see some of this progress could have been made after the early papers [15, 19, 16, 18] on what was then sometimes called the SY model. But the intervening developments on the connections between SYK, black holes, and holography were important in pointing to the correct direction, and Jan played a significant role in these. A possible holographic dual of the 2dYSYK model remains an interesting open question for future work, which we will not address here.
@@ 4 The ‘fan’: QPT without symmetry breaking
Indeed, the cuprate observations are the analog of those in the triangular lattice antiferromagnet KYbSe2 [63]. As discussed in Section 2 and Fig. 5, KYbSe2 displays a continuum at high energies which has been modeled in a theory of bosonic spinons [64, 63], along with weak antiferromagnetic order at low temperatures.
@@
Instead the observations of Robarts et al. [111] are naturally interpreted as spinons. Note the evolution of their observations with increasing doping from a sharp spin-wave spectrum in Figs. 8a,b to a broad continuum in Figs. 8c-f. This matches Fig. 9, which shows the evolution of the dynamic spin structure factor of the insulating square lattice antiferromagnet with increasing second neighbor exchange $J_{2}$, as computed by Ferrari and Becca [115] using fermionic spinons (see also the connection of spinons in the unfrustrated square lattice to the quasi-one-dimensional limit [116, 117]). There is also a correspondence in the energy scale, measured in units of the first neighbor exchange $J_{1}\approx 130$ meV [118].
@@ 5 Confinement crossover
It is worth mentioning here an important low $T$ difference between FL* and the holon metal which could be significant for transport experiments. As the holons are spinless, the area enclosed by a holon pocket is twice that of the hole pocket in FL* [146], and the holon pocket area is the same as that of the hole pocket in the ordered antiferromagnet. Very recent observations of the Yamaji effect [147] in the single layer cuprate HgBa2CuO4+δ at doping $p=0.1$ support the FL* theory of the pseudogap: there are 4 hole pockets of spin-1/2 fermions in FL*, and so each hole pocket occupies a fraction $p/8=0.0125$ of the Brillouin zone, to be compared to the observed value $\approx 1.3$% [147].
@@
The confinement crossover from the pseudogap metal to the stripy SDW state at low temperatures remains to be understood [157]. In the fermionic spinon approach, it involves a delicate interplay between the charge order associated with the condensation of $B$, and the confinement of the fermionic spinons via the SU(2) gauge field [141]. The bosonic spinon approach [75, 76, 77, 144] is an alternative route to understanding this crossover, and insights from numerical studies [158] should also be helpful.